# Hessian stability and convergence rates for entropic and Sinkhorn potentials via semiconcavity

### v1

@@ Abstract.
In this paper we determine quantitative stability bounds for the Hessian of entropic potentials, i.e., the dual solution to the entropic optimal transport problem. Up to authors’ knowledge this is the first work addressing this second-order quantitative stability estimate in general unbounded settings. Our proof strategy relies on semiconcavity properties of entropic potentials and on the representation of entropic transport plans as laws of forward and backward diffusion processes, known as Schrödinger bridges. Moreover, our approach allows to deduce a stochastic proof of quantitative stability entropic estimates and integrated gradient estimates as well. Finally, as a direct consequence of these stability bounds, we deduce exponential convergence rates for gradient and Hessian of Sinkhorn iterates along Sinkhorn’s algorithm, a problem that was still open in unbounded settings. Our rates have a polynomial dependence on the regularization parameter.
We warmly thank Alberto Chiarini and Giovanni Conforti for suggesting this problem and for many useful discussions. Both authors are associated to INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”) and the group GNAMPA. GG is supported by the PRIN project GRAFIA (CUP: E53D23005530006) and MUR Departement of Excellence Programme 2023-2027 MatMod@Tov (CUP: E83C23000330006). GG is thankful to Università Cattolica del Sacro Cuore for its hospitality and where part of this work has been carried out. LT acknowledges financial support by the INdAM-GNAMPA Project 2024 “Mancanza di regolarità e spazi non lisci: studio di autofunzioni e autovalori” (CUP: E53C23001670001).
@@ 1. Introduction
Given two probability measure $\rho,\,\mu\in\mathcal{P}(\mathbb{R}^{d})$ and a regularization parameter $T>0$, the Entropic Optimal Transport problem (EOT henceforth) reads as
@@
minimize $\text{minimize }\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{|x-y|^{2}}{2}\,\mathrm{d}\pi+T\,\mathscr{H}(\pi|\rho\otimes\mu)\text{ under the constraint }\pi\in\Pi(\rho,\mu)\;,$
@@
where $\mathscr{H}$ denotes the relative entropy functional (aka Kullback–Leibler divergence) and $\Pi(\rho,\mu)$ is the set of couplings of $\rho$ and $\mu$. This problem can be seen as an entropic regularization of the Optimal Transport (OT) problem, which indeed is recovered in the limit case $T=0$. For this reason, EOT has been widely studied in the last years and the solutions to its primal and dual formulation are respectively used as proxies for optimal transport plans and Brenier’s optimal transport map [42, 7, 46, 12]. Lastly, EOT is equivalent to a statistical mechanics problem, known as the Schrödinger problem, introduced in [52, 53] where E. Schrödinger was interested in the most likely evolution of a cloud of Brownian particles, conditionally to its initial and final distribution at time $s=0$ and $s=T$ respectively. Therefore EOT has a cutting-edge nature that lies at the interface between analysis and stochastics. Moreover, this problem has recently gained more popularity due to its use in machine learning and generative modeling applications [9, 57, 54], mainly due to the possibility of solving EOT via an iterative algorithm, known as Sinkhorn’s algorithm [55, 56] or Iterative Proportional Fitting Procedure, which can be used to quickly obtain approximate solutions for EOT [22] in a much easier and faster way, compared to standard OT solvers.
@@
In this article, we are interested in analyzing how changes in the marginals $\rho,\mu$ affect solutions to EOT. By relying on semiconcavity bounds and stochastic calculus, we are going to show below quantitative stability estimates for EOT potentials up to the second order, namely for their gradient and Hessian. To the best of our knowledge, this is the first work where second-order quantitative stability estimates are obtained. This is even more remarkable when compared with unregularized optimal transport, where higher-order quantitative stability estimates are more difficult to obtain, as in general potentials may lack regularity and the Ma–Trudinger–Wang condition [44] is imposed in order to ensure it; without this demanding assumption, only first-order quantitative stability bounds are available in general (see for instance the very recent [41, 37] and references therein). On the contrary, our main stability theorem is valid under fairly general assumptions, significantly weaker than the Ma–Trudinger–Wang condition, and since EOT is used as a proxy for OT, this highlights the importance of our result.
@@
In order to continue the exposition and state clearly our main contributions, let us collect a few basic facts about EOT and its solutions. First, let us recall that under mild assumptions on the marginals $\rho,\mu$ (see for instance [12, Proposition 2.2]), EOT admits a unique minimizer $\pi^{\mu}\in\Pi(\rho,\mu)$, referred to as the entropic plan (or Schrödinger plan), and there exist two functions $\varphi^{\mu}\in\mathrm{L}^{1}(\rho)$ and $\psi^{\mu}\in\mathrm{L}^{1}(\mu)$, called entropic potentials, such that
@@
Both the optimal plan $\pi^{\mu}$ and the entropic potentials $\varphi^{\mu},\psi^{\mu}$ depend on $T$ and on $\rho$, but for ease of notation we omit this dependence, as $T$ and $\rho$ will be kept fixed throughout the whole manuscript, whereas we are interested in stability bounds for changes in the second marginal in EOT. The couple $(\varphi^{\mu},\psi^{\mu})$ is unique up to constant translations $a\mapsto(\varphi^{\mu}+a,\psi^{\mu}-a)$ and it is characterized as solution to a system of equations. Indeed, if we suppose that the marginals admit densities of the form
@@
(1.1) $\begin{cases}\varphi^{\mu}=U_{\rho}+\log P_{T}\exp(-\psi^{\mu})\\ \psi^{\mu}=U_{\mu}+\log P_{T}\exp(-\varphi^{\mu})\;,\end{cases}$
@@
The structure of the Schrödinger system motivates the introduction of the ‘interpolated potentials’
@@
It is easily seen that they are solutions to the backward Hamilton-Jacobi-Bellman equation
@@
with final conditions $u_{T}=\varphi^{\mu}$ and $u_{T}=\psi^{\mu}$ respectively. Such a PDE enjoys a fundamental property of back-propagation of convexity (see Lemma A) and this has recently been employed in a stochastic analysis framework in order to prove convexity/concavity estimates for entropic potentials [20], providing an entropic version of the celebrated Caffarelli Theorem for Lipschitzianity of transport maps (see also [21, 31] for a non-stochastic proof). As shown in [15, 13], semiconcavity estimates play a pivotal role in establishing entropic quantitative stability results. In this work, we continue the research line started there, where semiconcavity was used for entropic stability of entropic plans and exponential convergence of Sinkhorn’s algorithm; here, we focus on quantitative stability bounds for gradient and Hessian of entropic potentials. For these reasons, let us introduce the notion of semiconcavity that we employ in our paper. We say that a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\Lambda$-semiconcave if for all $z,y\in\mathbb{R}^{d}$ we have
$f(z)-f(y)\leq\langle\nabla f(y),\,z-y\rangle+\frac{\Lambda}{2}\,|z-y|^{2}\,.$
As already observed in [15, 13], a crucial role is played by the semiconcavity of the function
@@
where $h\in\{\varphi^{\mu}_{0},\,\psi^{\mu}_{0}\}$ is a backpropagated entropic potential along HJB. We will denote with $\Lambda(h)$ a semiconcavity parameter [^1] To be more precise, in our examples and in the explicit computations we will fix a parameter $\Lambda\in\mathbb{R}$ such that (1) holds. We do not assume it to be the optimal parameter choice. of $g^{y}_{h}$ (uniformly in $y$).
@@
This first assumption is standard in EOT when considering its Schrödinger problem formulation and it guarantees the existence and uniqueness of optimal plan, entropic potentials as well as the validity of the stochastic representation via forward backward Schrödinger bridge processes, as described in Section 1.2 below. The second assumption is needed when introducing a different target marginal $\nu\in\mathcal{P}(\mathbb{R}^{d})$.
@@
Let us stress that, despite the finiteness of $\Lambda(\varphi^{\mu}_{0})$ in H 2 may seem as a condition on $\mu$, there exist sufficient conditions on $\rho$ that ensure its validity without any extra assumption on $\mu$. For instance, the compactness of the support of $\rho$ or the log-concavity of its Radon-Nikodým derivative, as discussed in Appendix A.
Under these assumptions, we will prove a general Hessian (and gradients) stability result which builds upon semiconcavity estimates for $\Lambda(\varphi^{\nu}_{0})$. In order to show its wide validity, we will further specialize these general estimates in two landmark examples: compactly supported and log-concave marginals. By building upon estimates obtained in [13] our quantitative stability estimates could be applied to weakly log-concave marginals or could be further specialized to the more regular Caffarelli’s setting (namely when the Hessian of marginals’ log-densities are both upper and lower bounded). For sake of exposition, we have omitted these two applications where the constants are less readable.
Our main stability result reads as follows.
@@
up to multiplicative constants that depend polynomially only on $\rho,\,\nu,\,T$ (and not on $\mu$). These constants are explicit and in particular, up to numerical universal constants, we have
  • if $\operatorname{supp}(\rho),\,\operatorname{supp}(\nu)\subseteq B_{R}(0)$ (for some radius big enough, i.e., $R^{2}\geq T$) then
@@
  • if both $\rho$ and $\nu$ are log-concave, i.e., their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\nu}\geq\alpha_{\nu}$ for some $\alpha_{\rho},\,\alpha_{\nu}>0$ (wlog such that $\alpha_{\rho}\vee\alpha_{\nu}<T^{-1}$), then
@@
In this paper whenever we write the $\mathrm{L}^{1}$-norm between matrices we are considering the $\mathrm{L}^{1}$-norm induced by the Hilbert-Schmidt norm of their difference, which is defined as
$\|A\|_{\mathrm{HS}}^{2}=\sum_{i,j}A_{i,j}^{2}\,.$
@@
Let us comment here on the dimensionality dependence in the previous bounds. More precisely, all our estimates are dimension-free up to being able to control the Hilbert-Schmidt norm of Hessian of backpropagated potentials $(\nabla^{2}\psi^{\nu}_{s})_{s\in[0,T)}$. In order to bound these last norms, we then rely on the known identity $\|\nabla^{2}\psi^{\nu}_{s}\|_{\mathrm{HS}}\leq\sqrt{d}\left\|\nabla^{2}\psi^{\nu}_{s}\right\|_{2}$ in (B.4) which allow us to efficiently bound this Hilbert-Schmidt in terms of the semiconcavity parameter $\Lambda(\psi^{\nu}_{0})$.
@@
Let us also comment on a hidden technical point and a first reason why the previous statement is “informal”. Rather than the differences $\nabla\varphi^{\nu}-\nabla\varphi^{\mu}$ and $\nabla^{2}\varphi^{\nu}-\nabla^{2}\varphi^{\mu}$, in Theorem 1.1 we control $\nabla(\varphi^{\nu}-\varphi^{\mu})$ and $\nabla^{2}(\varphi^{\nu}-\varphi^{\mu})$. Note indeed that no regularity assumptions are formulated on $\rho$, so that $\varphi^{\nu}$ and $\varphi^{\mu}$ may lack the required regularity, since they are not evolved along HJB. However, their difference $\varphi^{\nu}-\varphi^{\mu}$ is equal to $\psi^{\mu}_{0}-\psi^{\nu}_{0}$, which is instead the difference of two solutions to HJB, hence of two regular functions. Moreover, under some regularity assumption on $\rho$ (e.g. $\rho\in C^{2}(\mathbb{R}^{d})$), gradient and Hessian of $\varphi^{\mu},\varphi^{\nu}$ are in fact well defined. For this reason, in what follows we will always write $\nabla\varphi^{\nu}-\nabla\varphi^{\mu}$ and $\nabla^{2}\varphi^{\nu}-\nabla^{2}\varphi^{\mu}$, meaning $\nabla(\varphi^{\nu}-\varphi^{\mu})$ and $\nabla^{2}(\varphi^{\nu}-\varphi^{\mu})$ respectively whenever gradient/Hessian of the single potentials are not defined.
@@
For readers’ sake, we collect here the references within this article where our informal main result is stated and proven. The quantitative stability bound for gradients is proven in Theorem 3.1 whereas the Hessian stability bound is proven in Theorem 3.2, where the explicit constants are expressed in terms of $T$ and of the semiconcavity and geometric parameters of $\rho,\nu$. The above statement is informal also for a second reason: solely under H 1 and H 2, it is not clear whether these constants are finite, although we are able to show it and compute their asymptotics in our specialized setting. In particular, the compact setting bounds are given in Section 3.1 and Section 3.2, while for the log-concave setting in Section 3.1 and Section 3.2. These explicit bounds rely on the computations performed in Appendix A and Appendix B. Lastly, let us remark here that the above general result specifies also to different settings and our computations allow to derive explicit bounds also in the case when solely one marginal is compactly supported and the other one is log-concave. For sake of exposition we prefer not to insist on this, though these explicit bounds are a straightforward computation based on Appendix A and Appendix B and our general theorems.
Finally, it is clear that given the above the interested reader can deduce quantitative stability bounds when both marginals vary.

### v2

@@ Abstract.
In this paper we determine quantitative stability bounds for the Hessian of entropic potentials, i.e., the dual solution to the entropic optimal transport problem. To the authors’ knowledge this is the first work addressing this second-order quantitative stability estimate in general unbounded settings. Our proof strategy relies on semiconcavity properties of entropic potentials and on the representation of entropic transport plans as laws of forward and backward diffusion processes, known as Schrödinger bridges. Moreover, our approach allows to deduce a stochastic proof of quantitative stability estimates for entropic transport plans and for gradients of entropic potentials as well. Finally, as a direct consequence of these stability bounds, we deduce exponential convergence rates for gradient and Hessian of Sinkhorn iterates along Sinkhorn’s algorithm, a problem that was still open in unbounded settings. Our rates have a polynomial dependence on the regularization parameter.
Keywords: Entropic Optimal Transport; Hamilton-Jacobi-Bellman; Hessian stability; Schrödinger bridges; Sinkhorn’s algorithm
We warmly thank Alberto Chiarini and Giovanni Conforti for suggesting this problem and for many useful discussions. Both authors are associated to INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”) and the group GNAMPA. $G$ is supported by the PRIN project GRAFIA (CUP: E53D23005530006) and MUR Departement of Excellence Programme 2023-2027 MatMod@Tov (CUP: E83C23000330006). $G$ is thankful to Università Cattolica del Sacro Cuore for its hospitality and where part of this work has been carried out. LT acknowledges financial support by the INdAM-GNAMPA Project 2024 “Mancanza di regolarità e spazi non lisci: studio di autofunzioni e autovalori” (CUP: E53C23001670001).
@@ 1. Introduction
Given two probability measures $\rho,\,\mu\in\mathcal{P}(\mathbb{R}^{d})$ and a regularization parameter $T>0$, the Entropic Optimal Transport problem (EOT henceforth) reads as
@@
$\text{minimize }\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{|x-y|^{2}}{2}\,\mathrm{d}\pi+T\,\mathscr{H}(\pi|\rho\otimes\mu)\text{ under the constraint }\pi\in\Pi(\rho,\mu)\;,$
@@
where $\mathscr{H}$ denotes the relative entropy functional (aka Kullback–Leibler divergence) and $\Pi(\rho,\mu)$ is the set of couplings of $\rho$ and $\mu$. This problem can be seen as an entropic regularization of the Optimal Transport (OT) problem, which indeed is recovered in the limit case $T=0$. For this reason, EOT has been widely studied in the last years and the solutions to its primal and dual formulation are respectively used as proxies for optimal transport plans and Brenier’s optimal transport map [Mik04, BGN22, NW22, CCGT23]. Lastly, EOT is equivalent to a statistical mechanics problem, known as the Schrödinger problem, introduced in [Sch31, Sch32] where E. Schrödinger was interested in the most likely evolution of a cloud of Brownian particles, conditionally to its initial and final distribution at time $s=0$ and $s=T$ respectively. Therefore EOT has a cutting-edge nature that lies at the interface between analysis and stochastics. Moreover, this problem has recently gained more popularity due to its use in machine learning and generative modeling applications [BTHD21, WJX+21, SDBDD22], mainly due to the possibility of solving EOT via an iterative algorithm, known as Sinkhorn’s algorithm [Sin64, SK67] or Iterative Proportional Fitting Procedure, which can be used to quickly obtain approximate solutions for EOT [Cut13] in a much easier and faster way, compared to standard OT solvers.
@@
In this article, we are interested in analyzing how changes in the marginals $\rho,\mu$ affect solutions to EOT. By relying on semiconcavity bounds and stochastic calculus, we are going to show below quantitative stability estimates for EOT potentials up to the second order, namely for their gradient and Hessian. To the best of our knowledge, this is the first work where second-order quantitative stability estimates are obtained. This is even more remarkable when compared with unregularized optimal transport, where higher-order quantitative stability estimates are more difficult to obtain, as in general potentials may lack regularity and the Ma–Trudinger–Wang condition [MTW05] is imposed in order to ensure it; without this demanding assumption, only first-order quantitative stability bounds are available in general (see for instance the very recent [LM24, KLM25] and references therein). On the contrary, our main stability theorem is valid under fairly general assumptions, significantly weaker than the Ma–Trudinger–Wang condition, and since EOT is used as a proxy for OT, this highlights the importance of our result.
@@
In order to continue the exposition and state clearly our main contributions, let us collect a few basic facts about EOT and its solutions. First, let us recall that under mild assumptions on the marginals $\rho,\mu$ (see for instance [CCGT23, Proposition 2.2]), EOT admits a unique minimizer $\pi^{\mu}\in\Pi(\rho,\mu)$, referred to as the entropic plan (or Schrödinger plan), and there exist two functions $\varphi^{\mu}\in\mathrm{L}^{1}(\rho)$ and $\psi^{\mu}\in\mathrm{L}^{1}(\mu)$, called entropic potentials, such that
@@
Both the optimal plan $\pi^{\mu}$ and the entropic potentials $\varphi^{\mu},\psi^{\mu}$ depend on $T$ and on $\rho$, but for ease of notation we omit this dependence, as $T$ and $\rho$ will be kept fixed throughout the whole manuscript, whereas we are interested in stability bounds for changes in the second marginal in EOT. The pair $(\varphi^{\mu},\psi^{\mu})$ is unique up to constant translations $a\mapsto(\varphi^{\mu}+a,\psi^{\mu}-a)$ and it is characterized as solution to a system of equations. Indeed, if we suppose that the marginals admit densities of the form
@@
(1.1) $\varphi^{\mu}=U_{\rho}+\log P_{T}\exp(-\psi^{\mu})\,,\qquad\psi^{\mu}=U_{\mu}+\log P_{T}\exp(-\varphi^{\mu})\;,$
@@
The structure of the Schrödinger system motivates the introduction of the “interpolated potentials”
@@
It is easily seen that they are solutions to the backward Hamilton–Jacobi–Bellman equation
@@
with final conditions $u_{T}=\varphi^{\mu}$ and $u_{T}=\psi^{\mu}$ respectively. Such a PDE enjoys a fundamental property of backpropagation of convexity (see Appendix A in the Appendix) and this has recently been employed in a stochastic analysis framework in order to prove convexity/concavity estimates for entropic potentials in [Con24], providing an entropic version of the celebrated Caffarelli Theorem for Lipschitzianity of transport maps (see also [CP23, FGP20] for a non-stochastic proof). As shown in [CDG23, CCGT24], semiconcavity estimates play a pivotal role in establishing entropic quantitative stability results. In this work, we continue the research line started there, where semiconcavity was used for entropic stability of entropic plans and exponential convergence of Sinkhorn’s algorithm; here, we focus on quantitative stability bounds for gradient and Hessian of entropic potentials. For these reasons, let us introduce the notion of semiconcavity that we employ in our paper. We say that a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\Lambda$-semiconcave if for all $z,y\in\mathbb{R}^{d}$ we have
(1.2) $f(z)-f(y)\leq\langle\nabla f(y),\,z-y\rangle+\frac{\Lambda}{2}\,|z-y|^{2}\,.$
As already observed in [CDG23, CCGT24], a crucial role is played by the semiconcavity of the function
@@
where $h\in\{\varphi^{\mu}_{0},\,\psi^{\mu}_{0}\}$ is a backpropagated entropic potential along HJB. We will denote with $\Lambda(h)$ a semiconcavity parameter of $g^{y}_{h}$ (uniform in $y$). To be more precise, in our examples and in the explicit computations we will fix a parameter $\Lambda\in\mathbb{R}$ such that (1.2) holds. We do not assume it to be the optimal parameter choice.
@@
This first assumption is standard in EOT when considering its Schrödinger problem formulation and it guarantees the existence and uniqueness of optimal plan, entropic potentials as well as the validity of the stochastic representation via forward-backward Schrödinger bridge processes, as described in Section 1.2 below. The second assumption is needed when introducing a different marginal $\nu\in\mathcal{P}(\mathbb{R}^{d})$.
@@
Let us stress that, despite that the finiteness of $\Lambda(\varphi^{\mu}_{0})$ in H 2 may seem as a condition on $\mu$, there exist sufficient conditions on $\rho$ that ensure its validity without any extra assumption on $\mu$. For instance, the compactness of the support of $\rho$ or the log-concavity of its Radon–Nikodým derivative, as shown by the computations performed in the Appendix.
Under these assumptions, we will prove a general Hessian (and gradient) stability result which builds upon semiconcavity estimates for $\Lambda(\varphi^{\nu}_{0})$. In order to show its wide validity, we will further specialize these general estimates in two landmark examples: compactly supported and log-concave marginals. By building upon estimates obtained in [CCGT24] our quantitative stability estimates could be applied to weakly log-concave marginals or could be further specialized to the more regular Caffarelli’s setting (namely when the Hessian of marginals’ log-densities are both upper and lower bounded). For sake of exposition, we have omitted these two applications where the constants are less readable. In what follows, whenever we write that a constant depends polynomially on a measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$, we mean that it depends at most polynomially on the geometric parameters of $\rho$, such as the diameter of the support or the log-concavity parameter. Our main stability result reads as follows.
@@
up to multiplicative constants that depend polynomially only on $\rho,\,\nu,\,T$ (and not on $\mu$), and are explicit. Moreover, if we specify our result to the following settings we deduce that
  • If H 1 holds, $\mathscr{H}(\nu|\mathrm{Leb})<\infty$, $\operatorname{supp}(\rho),\,\operatorname{supp}(\nu)\subseteq B_{R}(0)$ (for some radius big enough, i.e., $R^{2}\geq T$) and either $\mu\ll\nu$ or $\operatorname{supp}(\mu)\subseteq B_{R}(0)$, then
@@
  • If H 1 holds, $\mathscr{H}(\nu|\mathrm{Leb})<\infty$, and both $\rho$ and $\nu$ are log-concave, i.e., their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\nu}\geq\alpha_{\nu}$ for some $\alpha_{\rho},\,\alpha_{\nu}>0$ (wlog such that $\alpha_{\rho}\vee\alpha_{\nu}<T^{-1}$), then
@@
In this paper, whenever we write the $\mathrm{L}^{1}$-norm of a matrix we are considering the $\mathrm{L}^{1}$-norm of its Hilbert–Schmidt norm, the latter being defined as $\|A\|_{\mathrm{HS}}^{2}=\sum_{i,j}A_{i,j}^{2}$. The presence of the dimension $d$ in these last second-order bounds comes from the Hilbert–Schmidt norm. Indeed, all our estimates are dimension-free up to being able to control the Hilbert–Schmidt norm of the Hessian of backpropagated potentials $(\nabla^{2}\psi^{\nu}_{s})_{s\in[0,T)}$. In order to bound these last norms, in the Appendix, we rely on the known identity $\|\nabla^{2}\psi^{\nu}_{s}\|_{\mathrm{HS}}\leq\sqrt{d}\left\|\nabla^{2}\psi^{\nu}_{s}\right\|_{2}$ which allows us to efficiently bound this Hilbert–Schmidt norm in terms of the semiconcavity parameter $\Lambda(\psi^{\nu}_{0})$.
Remark 2.
We have stated our general Wasserstein first- and second-order stability estimates under the absolute continuity assumption $\mu\ll\nu$. This restriction is mainly due to the stochastic control and entropic strategy we have employed in the proof of our main result. In particular, this is a necessary condition for the pivotal bound (2.6a) we prove for conditional Schrödinger bridges. Despite that, our Wasserstein stability estimates can be extended via a regularization argument to a more general setting where $\mu\not\ll\nu$, provided one is able to control the semiconcavity parameter alongside the regularization procedure. This requires a case-by-case discussion. In the compactly supported case we show that if we further assume $\operatorname{supp}(\mu)\subseteq B_{R}(0)$, then in H 2 we may drop $\mu\ll\nu$ (which is a stronger assumption). Moreover, we show via a heat kernel regularization that $\mu\ll\nu$ in H 2 can be completely dropped in the log-concave case. For the interested reader, these regularization procedures are performed in the proof of Corollaries B.1, B.2, B.3, and B.4 in the Appendix, where we further compute explicit rates and constants for these two specific settings.
@@
Let us also comment on a hidden technical point and a first reason why the previous statement is “informal”. Rather than the differences $\nabla\varphi^{\nu}-\nabla\varphi^{\mu}$ and $\nabla^{2}\varphi^{\nu}-\nabla^{2}\varphi^{\mu}$, in Theorem 1.1 we control $\nabla(\varphi^{\nu}-\varphi^{\mu})$ and $\nabla^{2}(\varphi^{\nu}-\varphi^{\mu})$. Note indeed that no regularity assumptions are formulated on $\rho$, so that $\varphi^{\nu},\varphi^{\mu}$ may lack the required regularity. However, $\varphi^{\nu}-\varphi^{\mu}=\psi^{\mu}_{0}-\psi^{\nu}_{0}$, which is instead the difference of two solutions to HJB, hence of two regular functions. Moreover, under some regularity assumption on $\rho$ (e.g. $\rho\in C^{2}(\mathbb{R}^{d})$), gradient and Hessian of $\varphi^{\mu},\varphi^{\nu}$ are in fact well defined.
@@
For the reader’s sake, we collect here the references within this article where our informal main result is stated and proven. The quantitative stability bound for gradients is proven in Theorem 3.1 whereas the Hessian stability bound is proven in Theorem 3.6, where the explicit constants are expressed in terms of $T$ and of the semiconcavity and geometric parameters of $\rho,\nu$. The above statement is informal also for a second reason: solely under H 1 and H 2, it is not clear whether these constants are finite, although we are able to show it and compute their asymptotics in our specialized setting. In particular, the compact setting and log-concave bounds are based on the explicit computations we perform in Section B.2, Section B.2, Section B.2 and Section B.2 in the Appendix.